OpenStudy (anonymous):

I need help I have no idea if i am doing this right. The question says solve the equation algebraically and graphically. V^2-5=8-2v^2 my final answer was v=√13. Is that right and how would I solve it graphically

OpenStudy (phi):

V^2-5=8-2v^2 I would combine the v^2 terms: add +2V^2 to both sides V^2 + 2V^2 - 5 =8 -2V^2 + 2V^2 on the right side, -2V^2 + 2V^2 is 0 (you have 2 V^2 and take way 2 V^2 to get no V^2 ) on the left you have 3 V^2 so you have 3 V^2 -5 =8 now add +5 to both sides. you get 3 V^2 = 13 divide both sides by 3 V^2 = 13/3 now take the square root of both sides V= \( \pm \sqrt{\frac{13}{3} } \)

OpenStudy (phi):

notice you get 2 answers, either plus or minus the square root of 13/3

OpenStudy (phi):

to solve it graphically, we could do this let V be x (because I always use x and y for graphs) 3V^2 = 13 becomes 3x^2=13 we could write this as 3x^2-13 =0 we can make this a function of x by saying y= 3x^2 -13 and solving for x when y =0 or we can graph the function, and look at the graph for where the function crosses the x-axis (i.e. where y=0) to graph this function, try some small numbers near zero: say x= -2, -1, 0 , 1 ,2 figure out the y values. plot the points, connect the dots